Circular No. 7730

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V445 PUPPIS A. A. Henden, Universities Space Research Association and U.S. Naval

Observatory, Flagstaff Station (USNOFS); R. M. Wagner, LBT Observatory; and S. G. Starrfield, Arizona State University, report that the following infrared photometry of V445 Pup was obtained in the pre-dawn sky with the USNOFS 1.55-m telescope(+ ASTROCAM): Sept. 29.61 UT, L′ = 5.74 ± 0.12 (cf. L = 2.63 ± 0.05 on Jan. 31.5; IAUC 7580); Oct. 4.54, K′ = 9.15 ± 0.02. However, V - and I-band CCD images obtained at large airmass in a brightening sky with the USNOFS 1.0-m telescope on Oct. 4.51 show no object brighter than V = 20 and I = 19.5 mag at the location of V445 Pup. Henden et al. remark: “The object is evidently shrouded in a thick and dense carbon dust shell, in view of the apparent over-abundance of carbon in the ejecta previously observed in infrared and optical spectra (cf. IAUC 7559 and 7717). Further observations are requested to follow the evolution of this unique object.”

1998 ST_{27}

L. A. M. Benner, Jet Propulsion Laboratory; M. C. Nolan, Arecibo Observatory; S. J. Ostro and J. D. Giorgini, Jet Propulsion Laboratory; and J. L. Margot, California Institute of Technology, write: “Arecibo (2380-MHz, 13-cm) delay-Doppler radar images obtained on Oct. 7 and 9 reveal that 1998 ST_{27} (MPEC 2001-S28) is a binary system. The images show separations up to at least 4 km between the components, whose sizes differ by a factor of at least three. Follow-up optical observations are desirable.”

COMETS C/1996 B5, C/1996 R4, C/2001 S2 (SOHO) Further to IAUC 7726, D. Hammer reports his measurements for ad-

ditional Kreutz sungrazing comets found by X. Leprette (C/1996 B5) and M. Oates on SOHO website C3 coronagraph images (and also C2 images for C/2001 S2). The reduced observations and orbital elements by B. G. Marsden are included on MPEC 2001-T15.

Comet  1996 UT  α_{2000}  δ_{2000}
C/1996 B5 Jan. 29.352  21^{h}01.^{m}4  −19^{o}25′
C/1996 R4 Sept.14.804  11 19.4  + 2 24
Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 S2 Sept.28.154  11^{h}58.^{m}1  − 3^{o}01′
2001 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green